VIBRATIONAL SPECTRA AND THEORETICAL CALCULATIONS FOR 2-METHYLCYCLOPENTANONE

\underline{ARNOLD DEL ROSARIO}, KEVIN MORRIS, JAAN LAANE, Department of Chemistry, Texas A&M University, College Station, TX 77843.

The liquid and vapor phase far-infrared, mid-infrared and Raman spectra of 2-methylcyclopentanone have been recorded and analyzed. Semi-empirical (MM3) and ab-initio (Gaussian 94) calculations have also been performed to help assign the ground state spectra and predict ground state conformations. The low frequency modes were of special interest: the ring bending (75 cm^{−1}), the ring twisting (251 cm^{−1}), and the methyl torsion (209 cm^{−1}). Preliminary fluorescence studies of the C=O inversion in the S_{1}(n,π^{∗}) state have been carried out.